Doc. Math. J. DMV 9

Section 1 Logic

In case of several authors, Invited Speakers are marked with a ^{∗}. Matthew Foreman: Generic Large Cardinals: New Axioms for

Mathematics? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 11

Greg Hjorth: When is an Equivalence Relation Classifiable? . . . . . . . II  23
Ludomir Newelski: Meager Forking and m-Independence . . . . . . . . . . II  33
Stevo Todorcevic: Basis Problems in Combinatorial Set Theory . . II  43

Documenta Mathematica · Extra Volume ICM 1998 · II ·

10

Documenta Mathematica · Extra Volume ICM 1998 · II ·